OpenStudy (anonymous):

If w is jointly proportional to u and v, and w =2 when u=4 and v=6, what is w when u=3 and v=8?

OpenStudy (anonymous):

\[\frac{2}{w}=\frac{4*6}{3*8}\]

OpenStudy (anonymous):

cross and time 2*3*8=w*4*6 48 =24w 48/24=w 2=w

OpenStudy (anonymous):

ty